# Universal Deformations and Inhomogeneities in Isotropic Cauchy Elasticity

### v1

@@ Abstract
For a given class of materials, universal deformations are those deformations that can be maintained in the absence of body forces and by applying solely boundary tractions. For inhomogeneous bodies, in addition to the universality constraints that determine the universal deformations, there are extra constraints on the form of the material inhomogeneities—universal inhomogeneity constraints. Those inhomogeneities compatible with the universal inhomogeneity constraints are called universal inhomogeneities. In a Cauchy elastic solid, stress at a given point and at an instance of time is a function of strain at that point and that exact moment in time, without any dependence on prior history. A Cauchy elastic solid does not necessarily have an energy function, i.e., Cauchy elastic solids are, in general, non-hyperelastic (or non-Green elastic). In this paper we characterize universal deformations in both compressible and incompressible inhomogeneous isotropic Cauchy elasticity. As Cauchy elasticity includes hyperelasticity, one expects the universal deformations of Cauchy elasticity to be a subset of those of hyperelasticity both in the compressible and incompressible cases. It is also expected that the universal inhomogeneity constraints to be more strict than those of hyperelasticity, and hence, the set of universal inhomogeneities to be smaller than that of hyperelasticity. We prove the unexpected result that the sets of universal deformations of isotropic Cauchy elasticity and isotropic hyperelasticity are identical, in both the compressible and incompressible cases. We also prove that their corresponding universal inhomogeneities are identical as well.
@@ 1 Introduction
For a given class of materials, universal deformations are those deformations that can be maintained in the absence of body forces and by applying only boundary tractions. Universal deformations have played a crucial role in nonlinear elasticity: i) They have had an important organizational role in the semi-inverse solutions in nonlinear elasticity (Tadmor et al., 2012; Goriely, 2017), and more recently in anelasticity (Kumar and Yavari, 2023) and viscoelasticity (Sadik and Yavari, 2024). ii) They offer guidance for designing experiments for determining the constitutive relations of a specific material (Rivlin and Saunders, 1951). iii) These exact solutions have been used as benchmark problems in computational mechanics (Chi et al., 2015; Shojaei and Yavari, 2018). iv) Universal deformations have been used in deriving effective properties for nonlinear composites (Hashin, 1985; Lopez-Pamies et al., 2012; Golgoon and Yavari, 2021).
@@
The study of universal deformations was extended to inhomogeneous anisotropic solids in (Yavari and Goriely, 2021; Yavari, 2021; Yavari and Goriely, 2023). [^1] Prior to our work, there had been some limited work on universal deformations in anisotropic solids (Ericksen and Rivlin, 1954). These comprehensive studies included both compressible and incompressible isotropic, transversely isotropic, orthotropic, and monoclinic solids. For these three classes of compressible anisotropic solids, it was shown that universal deformations are homogenous, and the material preferred directions are uniform. Additionally, for isotropic solids and each of the three classes of anisotropic solids, the corresponding universal inhomogeneities—these represent inhomogeneities in the energy function that are compatible with the universality constraints—were characterized. The corresponding universal inhomogeneities for each of the above six known families of universal deformations were determined for inhomogeneus isotropic and the three classes of inhomogeneous incompressible anisotropic solids.
@@
In linear elasticity, universal displacements are the counterparts to universal deformations (Truesdell, 1966; Gurtin, 1972; Yavari et al., 2020). Yavari et al. (2020) demonstrated the explicit dependence of universal displacements on the symmetry class of the material. Specifically, the larger the symmetry group, the larger the corresponding set of universal displacements. Therefore, isotropic solids have the largest set of universal displacements, while triclinic solids possess the smallest set of universal displacements. The investigation into universal displacements has also been extended to inhomogeneous solids (Yavari and Goriely, 2022b) and linear anelasticity (Yavari and Goriely, 2022a).
@@
A class of materials with internal constraints, significant in engineering applications, consists of materials reinforced with inextensible fibers (Adkins and Rivlin, 1955; Pipkin and Rogers, 1971; Pipkin, 1974, 1979). Despite their importance, there is a scarcity of literature on the universal deformations of fiber-reinforced solids. Beskos (1972) studied homogeneous compressible isotropic solids reinforced with inextensible fibers, investigating whether universal deformations of incompressible isotropic solids are universal for this class as well. Families $1$, $2$, $3$, and $4$ were specifically analyzed, revealing that certain subsets of these families are universal for specific fiber distributions. Interestingly, all these universal deformations are found to be homogeneous except for the shearing of a circular tube with circumferential fibers. Beatty (1978) further examined homogeneous compressible isotropic solids reinforced by a single family of inextensible fibers. He investigated the problem of finding all those fiber distributions for which homogenous deformations are universal. He proved that there are only three types of such fiber distributions, all characterized by straight fibers. In a recent study, Yavari (2023) studied universal displacements in compressible anisotropic linear elastic solids reinforced by a family of inextensible fibers. For each symmetry class, and under the assumption of a uniform distribution of straight fibers respecting the corresponding symmetry, the respective universal displacements were characterized. Interestingly, it was observed that, except for triclinic and cubic solids, the presence of inextensible fibers enlarges the set of universal displacements within the other five classes. [^2] It should be noted that a fiber-reinforced solid with straight fibers cannot be isotropic.
@@
In this paper, we extend the study of universal deformations and inhomogeneities to inhomogeneous compressible and incompressible isotropic Cauchy elasticity. Cauchy elastic solids may not necessarily have an energy function and include hyperelastic solids (Green elastic solids) as a special case. This suggests that Cauchy elasticity may have more stringent universality and universal inhomogeneity constraints compared to those of hyperelasticity. This, in turn, leads one to anticipate smaller sets of universal deformations and universal inhomogeneities for Cauchy elasticity compared to those of hyperelasticity. We prove the unexpected result that the universal deformations and inhomogeneities of Cauchy elasticity are identical to those of Green elasticity in both the compressible and incompressible cases.
@@ 2 Cauchy Elasticity
In Cauchy elasticity, stress at a point and a given moment in time is explicitly a function of strain at that point and that particular moment in time (Cauchy, 1828; Truesdell, 1952; Truesdell and Noll, 2004). However, an energy function does not necessarily exist. [^3] It is important to note that Cauchy elasticity does not encompass all elastic solids. In recent years, there has been some interest in implicit constitutive equations, e.g., constitutive equations of the form $\bm{\mathsf{f}}(\bm{\sigma},\mathbf{b})=\mathbf{0}$ (Rajagopal, 2003, 2007). Cauchy elasticity is a subset of this class of solids. In terms of the first Piola-Kirchhoff stress (Truesdell, 1952; Truesdell and Noll, 2004; Ogden, 1984)
@@
One can show that objectivity implies that (Truesdell and Noll, 2004)
@@
$\mathbf{S}=-p\,\mathbf{C}^{-1}+\Lambda_{0}\mathbf{G}^{\sharp}+\Lambda_{1}\mathbf{C}^{\sharp}\,,$ (2.4)
@@
where $\alpha=\alpha(X,I_{1},I_{2},I_{3})$, $\beta=\beta(X,I_{1},I_{2},I_{3})$, and $\gamma=\gamma(X,I_{1},I_{2},I_{3})$ are arbitrary functions. Similarly, for incompressible isotropic Cauchy elastic solids one has
@@
where $\beta=\beta(X,I_{1},I_{2})$ and $\gamma=\gamma(X,I_{1},I_{2})$ are arbitrary functions. In components, they read $\sigma^{ab}=\alpha\,g^{ab}+\beta\,b^{ab}+\gamma\,c^{ab}$ and $\sigma^{ab}=-p\,g^{ab}+\beta\,b^{ab}+\gamma\,c^{ab}$, respectively.
Green and Naghdi (1971) showed that Cauchy elasticity is consistent with the first and second laws of thermodynamics. They demonstrated that over a closed path in the space of strains the net work in a Cauchy elastic solid may not vanish. They explicitly demonstrated this in the case of a linear stress-strain relationship and observed that the lack of major symmetries is responsible for the nonzero work. [^4] The same thing was shown fifty years later in (Scheibner et al., 2020) when formulating the so-caled “odd” elasticity. These authors clearly were not aware of similar developments in the literature of nonlinear elasticity. We will critically examine “odd” elasticity in a separate communication. There is no consensus in the literature of nonlinear elasticity on Cauchy elasticity being a viable theory; while some dismiss it (Rivlin, 1986; Casey, 2005; Carroll, 2009; Rajagopal, 2011), others appear to accept it (Truesdell, 1952; Truesdell and Noll, 2004; Ogden, 1984; Kružík and Roubíček, 2019; Bordiga et al., 2022). Our motivation for studying Cauchy elasticity is because of its promise in describing the mechanics of active matter at large strains. Active matter may have access to external sources of energy, and hence, the net work done in a closed loop in the strain space may not vanish. The recent interest in the physics literature in the so-called “odd” elasticity (Scheibner et al., 2020; Fruchart et al., 2023), which is simply linearized non-hyperelastic Cauchy elasticity, is another motivation for revisiting Cauchy elasticity.
@@ 3 Universal Deformations in Compressible Isotropic Cauchy Elasticity
Let us consider a body made of compressible isotropic Cauchy elastic solids deforming in the Euclidean ambient space for which a single Cartesian coordinate chart $\{x^{a}\}$ can be chosen. Thus, $g_{ab}=\delta_{ab}$. In this coordinate chart the Cauchy stress has the following representation
$\sigma^{ab}=\alpha\,\delta^{ab}+\beta\,b^{ab}+\gamma\,c^{ab}\,.$ (3.1)
When there are no body forces present, the equilibrium equations in Cartesian coordinates read $\sigma^{ab}{}_{,b}=0$. Thus
$\sigma^{ab}{}_{,b}=\beta\,b^{ab}{}_{,b}+\gamma\,c^{ab}{}_{,b}+\alpha_{,b}\,\delta^{ab}+\beta_{,b}\,b^{ab}+\gamma_{,b}\,c^{ab}=0\,.$ (3.2)
@@
$\beta\,b^{ab}{}_{,b}+\gamma\,c^{ab}{}_{,b}$ (3.6)
$+I_{1,b}\,\delta^{ab}\,\alpha_{1}+I_{2,b}\,\delta^{ab}\,\alpha_{2}+I_{3,b}\,\delta^{ab}\,\alpha_{3}$
@@
$I_{1},I_{2},I_{3}~{}\text{are constants},\quad\text{and}\qquad b^{ab}{}_{,b}=c^{ab}{}_{,b}=0\,.$ (3.12)
Using these and compatibility equations, one can show that the universal deformations must be homogeneous (Ericksen, 1955). Thus, we have proved the following two results.
@@
We observe that the set of universal deformations of homogeneous compressible isotropic Cauchy elastic solids is identical to that of homogeneous compressible isotropic hyperelastic solids (Ericksen, 1955). Additionally, it is noteworthy that neither inhomogeneous compressible isotropic Cauchy elastic solids nor inhomogeneous compressible isotropic hyperelastic solids admit universal deformations (Yavari, 2021).
@@ 4 Universal Deformations in Incompressible Isotropic Cauchy Elasticity
These terms must be symmetric in $(a,b)$ for $A=1,2,3$ and $A\leq B=1,2,3$. The six terms highlighted in purple are identical to those of inhomogeneous isotropic hyperelasticity (Yavari, 2021).
@@
Also the universal inhomogeneity constraints of hyperelasticity are (Yavari, 2021):
@@
Next we show that the universal inhomogenietity constranits of Cauchy elasticity (4.23)-(4.23) are equivalent to those of hyperelasticity (4.40)-(4.46). We only need to check the terms in black, as the ones highlighted in purple are identical. Let us define $\mathsf{P}^{A}_{bn}=F^{-A}{}_{n}\,I_{1,b}+F^{-A}{}_{b}\,I_{1,n}$ and note that $\mathsf{P}^{A}_{nb}=\mathsf{P}^{A}_{bn}$. From (4.24) we know that

### v2

@@ Abstract
For a given class of materials, universal deformations are those deformations that can be maintained in the absence of body forces and by applying solely boundary tractions. For inhomogeneous bodies, in addition to the universality constraints that determine the universal deformations, there are extra constraints on the form of the material inhomogeneities—universal inhomogeneity constraints. Those inhomogeneities compatible with the universal inhomogeneity constraints are called universal inhomogeneities. In a Cauchy elastic solid, stress at a given point and at an instance of time is a function of strain at that point and that exact moment in time, without any dependence on prior history. A Cauchy elastic solid does not necessarily have an energy function, i.e., Cauchy elastic solids are, in general, non-hyperelastic (or non-Green elastic). In this paper we characterize universal deformations in both compressible and incompressible inhomogeneous isotropic Cauchy elasticity. As Cauchy elasticity includes hyperelasticity, one expects the universal deformations of Cauchy elasticity to be a subset of those of hyperelasticity both in the compressible and incompressible cases. It is also expected that the universal inhomogeneity constraints to be more stringent than those of hyperelasticity, and hence, the set of universal inhomogeneities to be smaller than that of hyperelasticity. We prove the somewhat unexpected result that the sets of universal deformations of isotropic Cauchy elasticity and isotropic hyperelasticity are identical, in both the compressible and incompressible cases. We also prove that their corresponding universal inhomogeneities are identical as well.
Dedicated to Professor Kaushik Bhattacharya on the occasion of his $60$th birthday.
@@ 1 Introduction
Within a given class of materials, universal deformations refer to those deformations that can be maintained in the absence of body forces and by applying only boundary tractions, for any member of the material class. Universal deformations do not depend on the particular material within the class. However, the boundary tractions necessary to sustain a universal deformation depend on the specific material. Universal deformations have played a crucial role in nonlinear elasticity and anelasticity (in the sense of Eckart (1948)): i) They have had an important organizational role in the semi-inverse solutions in nonlinear elasticity (Knowles, 1979; Polignone and Horgan, 1991; De Pascalis et al., 2009; Tadmor et al., 2012; Goriely, 2017), and more recently in anelasticity (Kumar and Yavari, 2023) and viscoelasticity (Sadik and Yavari, 2024). ii) They offer guidance for designing experiments for determining the constitutive relations of a specific material (Rivlin and Saunders, 1951; Saccomandi, 2001). [^1] The following quote from (Doyle and Ericksen, 1956, p. 89) explains it best: “From the standpoint of attempting to determine the form of $\Sigma$ for a given material by comparing general solutions with results of experiment, it appears that the solutions which are most useful are those which correspond to deformations which can be produced in every material of the type considered by the application of surface tractions only.” Here, $\Sigma$ is the energy function and by “general” they mean “universal”. iii) All the existing exact solutions of defects in nonlinear solids are related to universal deformations (Wesolowski and Seeger, 1968; Gairola, 1979; Zubov, 1997; Yavari and Goriely, 2012a, b, 2013a, 2014; Golgoon and Yavari, 2018b). iv) Universal deformations have been important for finding exact solutions for distributed finite eigenstrains in nonlinear solids and the nonlinear analogues of Eshelby’s inclusion problem (Yavari and Goriely, 2013b, 2015; Golgoon and Yavari, 2018a; Yavari, 2021b). v) These exact solutions have been used as benchmark problems in computational mechanics (Dragoni, 1996; Saccomandi, 2001; Chi et al., 2015; Shojaei and Yavari, 2018). vi) Universal deformations have been used in deriving effective properties for nonlinear composites (Hashin, 1985; Lopez-Pamies et al., 2012; Golgoon and Yavari, 2021).
@@
The study of universal deformations was extended to inhomogeneous anisotropic solids in (Yavari and Goriely, 2021; Yavari, 2021a; Yavari and Goriely, 2023). [^2] Prior to our work, there had been some limited work on universal deformations in anisotropic solids (Ericksen and Rivlin, 1954). These comprehensive studies included both compressible and incompressible isotropic, transversely isotropic, orthotropic, and monoclinic solids. For these three classes of compressible anisotropic solids, it was shown that universal deformations are homogenous, and the material preferred directions are uniform. Additionally, for isotropic solids and each of the three classes of anisotropic solids, the corresponding universal inhomogeneities—these represent inhomogeneities in the energy function that are compatible with the universality constraints—were characterized. The corresponding universal inhomogeneities for each of the above six known families of universal deformations were determined for inhomogeneus isotropic and the three classes of inhomogeneous incompressible anisotropic solids.
@@
In linear elasticity, universal displacements are the counterparts to universal deformations (Truesdell, 1966; Gurtin, 1972; Carroll, 1973; Yavari et al., 2020). Yavari et al. (2020) demonstrated the explicit dependence of universal displacements on the symmetry class of the material. Specifically, the larger the symmetry group, the larger the corresponding set of universal displacements. Therefore, isotropic solids have the largest set of universal displacements, while triclinic solids possess the smallest set of universal displacements. The investigation into universal displacements has also been extended to inhomogeneous solids (Yavari and Goriely, 2022b) and linear anelasticity (Yavari and Goriely, 2022a).
@@
A class of materials with internal constraints, significant in engineering applications, consists of materials reinforced with inextensible fibers (Adkins and Rivlin, 1955; Pipkin and Rogers, 1971; Pipkin, 1974, 1979). Despite their importance, there is a scarcity of literature on the universal deformations of fiber-reinforced solids. Beskos (1972) studied homogeneous compressible isotropic solids reinforced with inextensible fibers, investigating whether universal deformations of incompressible isotropic solids are universal for this class as well. Families $1$, $2$, $3$, and $4$ were specifically analyzed, revealing that certain subsets of these families are universal for specific fiber distributions. Interestingly, all these universal deformations are found to be homogeneous except for the shearing of a circular tube with circumferential fibers. Beatty (1978) further examined homogeneous compressible isotropic solids reinforced by a single family of inextensible fibers. He investigated the problem of finding all those fiber distributions for which homogenous deformations are universal. He proved that there are only three types of such fiber distributions, all characterized by straight fibers. In a recent study, Yavari (2023) studied universal displacements in compressible anisotropic linear elastic solids reinforced by a single family of inextensible fibers. For each symmetry class, and under the assumption of a uniform distribution of straight fibers respecting the corresponding symmetry, the respective universal displacements were characterized. Interestingly, it was observed that, except for triclinic and cubic solids, the presence of inextensible fibers enlarges the set of universal displacements within the other five classes. [^3] It should be noted that a fiber-reinforced solid with straight fibers cannot be isotropic.
@@
In this paper, we extend the study of universal deformations and inhomogeneities to inhomogeneous compressible and incompressible isotropic Cauchy elasticity. Cauchy elastic solids may not necessarily have an energy function and include hyperelastic solids (Green elastic solids) as a special case. This suggests that Cauchy elasticity may have more stringent universality and universal inhomogeneity constraints compared to those of hyperelasticity. This, in turn, leads one to anticipate smaller sets of universal deformations and universal inhomogeneities for Cauchy elasticity compared to those of hyperelasticity. We prove the somewhat unexpected result that the universal deformations and inhomogeneities of Cauchy elasticity are identical to those of Green elasticity in both the compressible and incompressible cases.
@@ 2 Cauchy Elasticity
In Cauchy elasticity, stress at a point and a given moment in time is explicitly a function of strain at that point and that particular moment in time (Cauchy, 1828; Truesdell, 1952; Truesdell and Noll, 2004). However, an energy function does not necessarily exist. [^4] It is important to note that Cauchy elasticity does not encompass all elastic solids. In recent years, there has been some interest in implicit constitutive equations, e.g., constitutive equations of the form $\bm{\mathsf{f}}(\bm{\sigma},\mathbf{b})=\mathbf{0}$ (Morgan, 1966; Rajagopal, 2003, 2007). Cauchy elasticity is a subset of this class of solids. In terms of the first Piola-Kirchhoff stress (Truesdell, 1952; Truesdell and Noll, 2004; Ogden, 1984)
@@
One can show that objectivity implies that the second Piola-Kirchhoff stress has the following functional form (Truesdell and Noll, 2004)
@@
$\mathbf{S}=-p\,\mathbf{C}^{-\sharp}+\Lambda_{0}\mathbf{G}^{\sharp}+\Lambda_{1}\mathbf{C}^{\sharp}\,,$ (2.4)
@@
where $\alpha=\alpha(X,I_{1},I_{2},I_{3})$, $\beta=\beta(X,I_{1},I_{2},I_{3})$, and $\gamma=\gamma(X,I_{1},I_{2},I_{3})$ are arbitrary response functions. Similarly, for incompressible isotropic Cauchy elastic solids one has
@@
where $\beta=\beta(X,I_{1},I_{2})$ and $\gamma=\gamma(X,I_{1},I_{2})$ are arbitrary response functions. In components, they read $\sigma^{ab}=\alpha\,g^{ab}+\beta\,b^{ab}+\gamma\,c^{ab}$ and $\sigma^{ab}=-p\,g^{ab}+\beta\,b^{ab}+\gamma\,c^{ab}$, respectively.
Green and Naghdi (1971) showed that Cauchy elasticity is consistent with the first and second laws of thermodynamics. They demonstrated that over a closed path in the space of strains the net work in a Cauchy elastic solid may not vanish. They explicitly demonstrated this in the case of a linear stress-strain relationship and observed that the lack of major symmetries is responsible for the nonzero work. [^5] The same thing was shown fifty years later in (Scheibner et al., 2020) when formulating the so-called “odd” elasticity. These authors clearly were not aware of similar developments in the literature of nonlinear elasticity. There is no consensus in the literature of nonlinear elasticity on Cauchy elasticity being a viable theory; while some dismiss it (Rivlin, 1986; Casey, 2005; Carroll, 2009; Rajagopal, 2011), others appear to accept it (Truesdell, 1952; Ericksen, 1956; Truesdell and Noll, 2004; Ogden, 1984; Kružík and Roubíček, 2019; Bordiga et al., 2022). Our motivation for studying Cauchy elasticity is because of its promise in describing the mechanics of active matter at large strains. Active matter may have access to external sources of energy, and hence, the net work done in a closed loop in the strain space may not vanish. The recent interest in the physics literature in the so-called “odd” elasticity (Scheibner et al., 2020; Fruchart et al., 2023), which is simply linearized non-hyperelastic Cauchy elasticity, is another motivation for revisiting Cauchy elasticity.
@@ 3 Universal Deformations in Compressible Isotropic Cauchy Elasticity
Let us consider a compressible isotropic Cauchy elastic body deforming in the Euclidean ambient space. In a local coordinate chart $\{x^{a}\}$, which may be curvilinear, the Cauchy stress has the following representation
$\sigma^{ab}=\alpha\,g^{ab}+\beta\,b^{ab}+\gamma\,c^{ab}\,.$ (3.1)
[…]
$\sigma^{ab}{}_{|b}=\beta\,b^{ab}{}_{|b}+\gamma\,c^{ab}{}_{|b}+\alpha_{,b}\,g^{ab}+\beta_{,b}\,b^{ab}+\gamma_{,b}\,c^{ab}=0\,.$ (3.2)
@@
$\beta\,b^{ab}{}_{|b}+\gamma\,c^{ab}{}_{|b}$ (3.6)
$+I_{1,b}\,g^{ab}\,\alpha_{1}+I_{2,b}\,g^{ab}\,\alpha_{2}+I_{3,b}\,g^{ab}\,\alpha_{3}$
@@
$I_{1},I_{2},I_{3}~{}\text{are constant},\quad\text{and}\qquad b^{ab}{}_{|b}=c^{ab}{}_{|b}=0\,.$ (3.12)
Using these and the compatibility equations, one can show that the universal deformations must be homogeneous (Ericksen, 1955). Thus, we have proved the following two results.
@@
We observe that the set of universal deformations of homogeneous compressible isotropic Cauchy elastic solids is identical to that of homogeneous compressible isotropic hyperelastic solids (Ericksen, 1955). Additionally, it is noteworthy that neither inhomogeneous compressible isotropic Cauchy elastic solids nor inhomogeneous compressible isotropic hyperelastic solids admit universal deformations (Yavari, 2021a).
@@ 4 Universal Deformations in Incompressible Isotropic Cauchy Elasticity
These terms must be symmetric in $(a,b)$ for $A=1,2,3$ and $A\leq B=1,2,3$. The six terms highlighted in purple are identical to those of inhomogeneous isotropic hyperelasticity (Yavari, 2021a).
@@
Also the universal inhomogeneity constraints of hyperelasticity are (Yavari, 2021a):
@@
Next we show that the universal inhomogeneity constraints of Cauchy elasticity (4.23)-(4.30) are equivalent to those of hyperelasticity (4.40)-(4.46). We only need to check the terms in black, as the ones highlighted in purple are identical. Let us define $\mathsf{P}^{A}_{bn}=F^{-A}{}_{n}\,I_{1,b}+F^{-A}{}_{b}\,I_{1,n}$ and note that $\mathsf{P}^{A}_{nb}=\mathsf{P}^{A}_{bn}$. From (4.24) we know that